OpenStudy (anonymous):

A set of numbers is shown below: {5.4, 6.2, 7.4, 9.2, 15.1} The number __ from the set will make the equation x + 3 = 8.4 true.

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's say we pick a random number 15.1 from the set replace x with 15.1 to see if x+4 = 8.4 is true x+4 = 8.4 15.1+4 = 8.4 19.1 = 8.4 since 19.1 = 8.4 is false, this means x = 15.1 is NOT the solution to x + 4 = 8.4

jimthompson5910 (jim_thompson5910):

you are looking for values that make x + 4 = 8.4 true

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x + 3 = 8.4 true.

jimthompson5910 (jim_thompson5910):

try out other numbers from that set

OpenStudy (anonymous):

9.2

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Jack_Prism

jimthompson5910 (jim_thompson5910):

so 9.2 plus 3 is equal to 8.4 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ohhhhhhh thx